Hello.
I have the question of Given the following functions, find each of the values:
f(x)=−49
g(x)=x+7
(f+g)(-1)=
(f-g)(-2)=
(f*g)(-5)=
(f/g)(2)=

In order, I answered the following:
-53.9 (incorrect)
-39 (incorrect)
-49x+1715 (Incorrect)
(-49)/(x+7)(incorrect).
Any assistance on what I did wrong is appreciated. Thanks!

Answer by rothauserc(4718)

:
add the two functions
:
-49 + x + 7 = x - 42
:
(f+g)(-1) = -1 - 42 = -43
:
subtract the two functions
:
-49 - (x + 7) = -x - 56
:
(f-g)(-2) = -(-2) - 56 = -54
:
multiply the two functions
:
-49 * (x+7) = -49x -343
:
(f*g)(-5) = (-49 * -5) -343 = -98
:
divide the two functions
:
-49 / (x+7)
:
(f/g)(2) = -49 / (2 + 7) = -49 / 9 = -5.44444 approximately -5.4
: